CRY+VIS Interface Meeting
Date: 2021/10/8 13:00-13:30
Participants: Takahashi, Sato, Ikeda, Uchiyama, Kimura, Aso, Ushiba, Hirata, Akutsu
- IX: Health check. Drift of the geophone signals. IP control with the new LVDT driver.
- EX: Health check. Replacement of the LVDT driver for the IP/F0.
- IY: Recovery of the yaw picomotor in BF.
- EY: Adjustment of the IP position. Repair of the insulator was scheduled on 12-13.
- TM height setting: ITMY are ETMY are too low.
- Commissioning in air: Health check.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Under the bidding process.
- Heater control: Installed the heaters to BS and SRs.
Date: 2021/10/8 14:00-14:25
Participants: Takahashi, Sato, Ushiba, Fabian, Terrence, Tanaka, Aso, Sugiyama, Uchiyama, Hirata
- Recovered the yaw picomotor of BF in IY.
- Adjusted the position of IP in EY.
- Replaced the LVDT driver for IP in EX.
- Assembled the geophone pod.
Type-B and Type-Bp various
- On Tuesday, Hirata-san
- Connected Pico cable for PRM.
- Connected in-air cable for BS IP Yaw Stepper, and confirm limit switch works.
- Checked Ikeda-san's temperature controller cable inside the box with Yano-san and Ikeda-san.
- Closed SR2 chamber with Yano-san.
I uploaded to JGWDoc site a document about the V1 OSEM situation (JGW-T2113352-v1). I suggest there we should extend the measuring range to at least 1.3 mm. We should aim for a slightly larger range like 1.4 mm if possible.
- On Wednesday, we moved V1 OSEM from 280 to 3380 cnt. Our aim was 2000 counts but it's hard to aim. The idea was to have the OSEM as close as possible in range after correcting the pitch produced by the following earthquake.
- On Thursday we fixed a couple of bad connections in BF Damper LVDTs.
There was an earthquake yesterday and PR3 pitch changed by -10.2 mrad. Likely, close to where it was before the earthquake on the 19th of september.
- We corrected the pitch using the IM-P picomotor. We oplev beam is almost in the QPD centre.
- The V1 OSEM went from 1081 cnt to 10,700 cnt.
- We measures IM-P transfer shows the OSEM is not completely out of range!
- On Monday we should move the OSEM by 1000 counts or so.
- Installed the geophone into the OMC chamber.
- Type-A team: EY (Height adjustment, IP FR, H4 geophone), IY (Alignment, LVDT driver)
- Type-B team: PR3 OSEM, Cable repair, Geophone noise level
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
(Date in Kamioka)
The week from 10/11
- Takahashi: 10/12-15
- Sato: 10/12-15
- Hirata: 10/11-15
- Yano: 10/11-15
The week from 10/18
- Hirata: 10/19-22
- Yano: 10/18-22